OpenStudy (anonymous):

Solve the permutation. 11 P 9 / 11 P 8

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

do you know what is permutation?

OpenStudy (anonymous):

Yeah, I think. like for the 11 P 9, you would do: 11*10*9 right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[11\text{ P }9=\frac{11!}{2!}\] yes?

OpenStudy (anonymous):

Yeah, I understand so far.

OpenStudy (anonymous):

isn't that factorial?

OpenStudy (anonymous):

yes. and also \[11\text{ P }8=\frac{11!}{3!}\] so when you divide, invert and multiply

OpenStudy (anonymous):

11! = 11*10*9*8*7*6*5*4*3*2*1?

OpenStudy (anonymous):

\[nPk=\frac{n!}{(n-k)!}\] that's formula for permutations

OpenStudy (anonymous):

so you get a huge cancellation are are left with very little

OpenStudy (anonymous):

thomas.a wrote the formula, i used the numbers. you are dividing one by the other. invert and multiply cancel the 11! and then see what you get

OpenStudy (anonymous):

11! / 2! = 19958400?

OpenStudy (anonymous):

i know that you did but i wrote in general form and it's explained how you got denominator

OpenStudy (anonymous):

you should do like satellite advised

OpenStudy (anonymous):

k

OpenStudy (anonymous):

whenever you are dealing with factorials, cancel first, multiply last

OpenStudy (anonymous):

Cancel 11!?

OpenStudy (anonymous):

yes. there is one in the numerator and one in the denominator. out they go

OpenStudy (anonymous):

okay so 11! is no longer there. Just 2!, and then 3!?

OpenStudy (anonymous):

\[\frac{11!}{2!}\div \frac{11!}{3!}=\frac{11!}{2!}\times \frac{3!}{11!}\]

OpenStudy (anonymous):

and just to confirm, 11! would mean you do this math?: 11*10*9*8*7*6*5*4*3*2*1

OpenStudy (anonymous):

that is the meaning of 11! yes

OpenStudy (anonymous):

but in this case you do no such thing. just cancel them to get \[\frac{3!}{2!}=\frac{3\times 2}{2}=3\]

OpenStudy (anonymous):

So all you do to find this answer is find 11! * 3!, and 2! * 11, then divide the two large numbers?

OpenStudy (anonymous):

cancel cancel cancel. \[\frac{11!}{11!}=1\]

OpenStudy (anonymous):

okay i got it now. thanks.